You are given the following information about an economy. C=100+0.7Yd, I=100, G=500 and tax rate is 20%.
(i)Find the equilibrium level of income.
(ii) How much increase in income will take place if government expenditure on goods and services increased by k70?
(iii) Suppose the marginal propensity to save is 0.6, the marginal tax rate is 0.5 and the marginal propensity to import is 0.8. What will be the size of multiplier?
1
Expert's answer
2020-11-23T05:39:32-0500
"Yd=(1-t)Y"

"Y=100+0.7\times0.8Y+100+500"


"0.44Y=700"


"Y=1591"

ii)


"0.44Y=770"


"Y=1750"

iii)


"MPS=0.6"


"m'=0.8"


"t=0.5"

"MULT=\frac {1}{1-(MPS-m')(1-t)}=\frac {1}{1+0.2\times0.5}=0.9"